# Bounding Boxes and Probabilistic Graphical Models: Video Anomaly Detection Simplified

### v1

@@ Appendix G.0.4 Average Inference Time (sec)
Appendix H Bounding Box - Grid Cell Intersections
(a) Whole Box - Frame-level AUC: $55.54$, RBDC: $28.12$, TBDC: $44.15$
(b) Bottom Border - Frame-level AUC: $74.00$, RBDC: $60.63$, TBDC: $98.59$
@@ Appendix I.1 Explainability
The example result explainability visualization shown in Figure 4 of the main paper presents an explanation of the anomaly score that was extracted from one of the three affected grid cells. To show the complete picture for the object, we present the explanations for the remaining two cells in Figure 6. Similarly to the observations made in Section 4 in the main paper the results presented in this figure support the reason for the low probability score given by the cell analyzed in Figure 4: Due to the detection of a very odd bounding box size, velocity and direction the running person was flagged as an abnormal object in the frame. The only difference this time is that the most likely value for the velocity direction in cell 106 is ’none’ instead of ’E’.
Figure 6: Supplement to Figure 4 in Main Paper - Result Explanation for the Remaining Affected Cells in the Frame

### v2

@@ Appendix G.0.4 Average Inference Time (sec)
Appendix H Bounding Box - Grid Cell Intersections
(a) Whole Box - Frame-level AUC: $55.54$, RBDC: $28.12$, TBDC: $44.15$
(b) Bottom Border - Frame-level AUC: $74.00$, RBDC: $60.63$, TBDC: $98.59$
@@ Appendix I.1 Explainability
In the following, we present explanations for example detections of anomalous events for all three benchmark datasets. The examples were chosen to cover the widest range of possible anomalies. For transparency reasons, we also cover an example in which our method failed to detect the correct anomaly in the video.
Appendix I.1.1 CUHK Avenue
In contrast to the example visualization in the main paper, we present a per cell explanation of the anomaly score for a sample image taken from the test video $03$ of CUHK Avenue. Here, a person is moving at an unusual speed through the scene which is considered anomalous in the given context. This example represents a motion-dependent anomaly.
Figure 6: Temporal Anomaly: Explainability dashboards for test video $03$ of the benchmark dataset CUHK Avenue of all affected cells in which the given object appears.
From all three dashboards it can be deducted that apart from the velocity direction ("SW"), it is primarily the size ("x-large") and velocity ("lightning fast") of the given bounding box which are contributing most to the object’s low probability. The reason for this is that first, the measured values within those two categories do not correspond to the most likely ones, i.e., the ones occupying the largest fraction of the respective chart, and that second, they in fact represent the least likely values among all others that were encountered during training. With the bottom border of the object’s bounding box spanning three cells in total we thus present three explainability dashboards in Figure 6.
Appendix I.1.2 StreetScene
In the example chosen from test video $31$ from StreetScene on the other hand, the present anomaly depicts a cyclist driving on the car lane of a street. Here, the anomaly is primarily related to its spatial location. Since under normal circumstances, i.e., during training, the most common class of objects detected in this area are vehicles (here, ’car’ occupies $97\%$ of the chart area), the detected class ’person’ contributes most to the low probability score of the object. The same can be concluded about the observed value of the RV Velocity. In this case, the most probable velocity would be ’normal’ while the actual detected value corresponds to ’very fast’. Similarly to the example in the main paper, here we present the average probability distributions over all affected cells.
Figure 7: Spatial Anomaly: Explainability dashboard for test video $31$ of the benchmark dataset StreetScene of all affected cells in which the given object appears. A success case.
A few frames earlier in the same test video we chose to present the explainability charts for another object depicting a person sitting on the sidewalk which is also considered anomalous behavior. For transparency reasons, this particular example is meant to show in which cases our method fails to correctly identify anomalies, and shown in Figure 8. As shown by the visualization, the probability of this particular object appearing in this location of the frame amounts to over $90\%$, and is thus very likely according to our discrete BN. The observed values of nearly all RVs always correspond to the most likely ones. The only two exceptions are object movement ’Velocity’ and ’Direction’. In both cases, the observed values are equal to the second-most likely ones, i.e., ’idle’ and ’none’, respectively. The most likely ones in this particular case, ’undefined’, denote edge scenarios in which the object was not found in the preceding frame, leaving both, movement velocity and direction undefined.
Figure 8: Spatio-Temporal Anomaly: Explainability dashboard for test video $31$ of the benchmark dataset StreetScene of all affected cells in which the given object appears. A failure case.
The most apparent reason for this false negative is the division of the frame into grid cells of a size that covers the entire width of the sidewalk. This implies that all pedestrians that pass by/shortly stop in this area during training will contribute significantly to making a person that sits there for a long time appear normal.